OpenStudy (anonymous):

A steel sphere is dropped vertically onto a horizontal metal plate. The sphere hits the plate with a speed u, leaves it at speed v, and rebounds vertically to half its original height. Which expression gives the value of v/u?

OpenStudy (anonymous):

a) 1/2√ b)1−1/2√ c) 1/2^2 d) 1/2

OpenStudy (anonymous):

compare KE and PE from stage1 to KE and PE of stage 2\[ \frac{ \frac{ 1 }{ 2 }m u ^{2} }{ \frac{ 1 }{ 2 }m v^2 } = \frac{ mgh }{ mg \frac{ h }{ 2 } }\]

OpenStudy (anonymous):

Then what?

OpenStudy (anonymous):

can u please show me the steps

OpenStudy (anonymous):

I remember seeing someone answer this. I think the answer is 1/sqrt(2)?

OpenStudy (anonymous):

Correct